OpenStudy (anonymous):

Can someone walk me throw the steps into solving this system of equations?? 3x + y = 6 x - 2y = 2

OpenStudy (smurfy14):

add 2y to each side for second equation and get x=2y+2 then plug that in to first equation: 3(2y+2)+y=6 then distribute 6y+6+y=6 7y+6=6 7y=0 y=0

OpenStudy (anonymous):

thank you